Many metals react with halogens to give metal halides. For example,
2 Al (s) + 3 Cl2(g) → 2 AlCl3 (s)
If you begin with 13.5 g of aluminum
1
Expert's answer
2018-04-07T08:24:11-0400
To find mass of aluminum chloride, we should find moles of aluminum. Moles of aluminum chloride AlCl3 are equal moles of aluminum, because they both have the same coefficient it the reaction:
2Al(s) + 3Cl2(g) → 2AlCl3(s)
n(AlCl_3 )=n(Al)=(m(Al))/(M(Al))=(13.5 g)/(26.98 g/mol)=0.5 mol
m(AlCl_3 )=n(AlCl_3 )×M(AlCl_3 )=0.5 mol×133.34 g/mol=66.7 g
Answer:
If you begin with 13.5 g of aluminum, you can get 66.7 g of AlCl3.